# Taylor’s hypothesis and its impact on flux measurements in a forest clearcut

### v1

@@ Appendix C Spectral conversion
$f$ $=\frac{U_{c}\kappa}{2\pi{(1+\kappa z)}^{4/5}}$ (21)
$df$ $=\frac{U_{c}(5+\kappa z)}{10\pi{(1+\kappa z)}^{9/5}}d\kappa,$
where $U_{C}$ is the bulk convective speed and $z$ is the measurement height (3.1 m). By substituting them in Eq. 20, we obtain the following spectrum in spatial domain,
$S_{ww}(\kappa)=\frac{aU_{C}{(2\pi)}^{5/3}(5+\kappa z)}{10\pi{(1+\kappa z)}^{7/15}{(2\pi{(1+\kappa z)}^{4/5}+b\kappa U_{C})}^{5/3}}.$ (22)
For the same 30-min run, we show this parametric spectrum in Fig. 11b. In contrast to Fig. 11a, Eq. 22 modifies the $-5/3$ spectral slope with a one equal to $-1.25$. Therefore, the scale-dependent convective speeds do not yield the desired result for the vertical velocity.
This conclusion is even supported by the real data. For illustration purposes, corresponding to the along-wind cases, we plot the ensemble-averaged vertical velocity spectra in Figs. 11c–d. In Fig. 11c, the frequencies are converted to wavenumbers through a linear transformation, i.e. using a constant wind speed ($U$). On the other hand, in Fig. 11d, the conversion is carried out with scale-dependent convective speeds ($U_{C}(\kappa z)$). Since the wavenumbers and spectral densities are normalized in Figs. 11c–d, their ranges are different than Figs. 11a–b. Similar to Fig. 11b, in Fig. 11d, the conversion with scale-dependent convective speeds impacts the $-5/3$ spectral slope and replaces it with $-1.25$. Instead, one recovers a $-5/3$ scaling with a constant convective speed (Fig. 11c). Contrary to $w$, the scalars show a clear $-5/3$ spectral slope when the scale-dependent convective speeds are used (Fig. 7b). Thus, in terms of how fast the eddies move, dissimilarities exist between the scalars and vertical velocities. At present, this is a hypothesis and needs to be confirmed through large eddy simulations.

### v2

@@ Acknowledgements.
Figure 11: (a) The estimations of scale-dependent convection speeds are carried out for the randomly-shuffled DTS data. In accordance with Eq. 5, for the randomly-shuffled DTS data, scale-dependent convection speeds follow a 1-1 slope (shown as blue dash-dotted line). (b) The histogram ($N(\alpha)$) of the scale-dependent convection speed exponents ($\alpha$) is shown for 100 random bootstrapped samples with replacement, corresponding to the DTS temperature signal from Section 6. The vertical line indicates the exponent 0.8-0.8.
For our DTS set up, three types of noises could have impact on the exponent of the scale-dependent convection speeds. The first scenario is a pure white noise. peltola2021suitability showed that at smaller scales of the flow, the DTS measurements could be affected by the presence of a white noise. To test that part, we carried out a randomly-shuffled exercise on the DTS temperature data. At each instant of time, we randomly shuffled the spatial DTS data, thereby destroying any coherence between them. For randomly-shuffled DTS data, the phase differences between the spatial temperature signals at time instants $t$ and $t+\Delta t$ would be equal to $\pi/2$ at all the wavenumbers since no coherence is expected between them. For such scenarios, Eq. 5 suggests that the $U_{C}(\kappa)$ values would vary as ${(\kappa)}^{-1}$, since both $\varphi$ and $\Delta t$ remain constant irrespective of different $\kappa$. This is indeed the case as shown in Fig. 11a. As opposed to the real signals, for the randomly-shuffled samples, one can clearly see that the $U_{C}(\kappa z)$ values vary as ${(\kappa z)}^{-1}$ instead of ${(\kappa z)}^{-0.8}$. However, since the 1-1 exponent is not very different from 0.8-0.8, we carried out an additional statistical significance test. Corresponding to every section of the DTS cable, we generated 100 random bootstrapped samples with replacement of the original temperature data. Subsequently, for each of these bootstrapped samples, to compute the power-law exponent ($\alpha$), a linear regression was carried out between $\log[U_{C}(\kappa z)/U_{C}]$ and $\log[\kappa z]$ at wavenumbers $\kappa z>1$. To illustrate the outcome, in Fig. 11b, we show a histogram of these exponents ($N(\alpha)$) for section-6 of the DTS cable (i.e. perpendicular-wind direction). It is clear that these $\alpha$ values cluster around 1-1 with 0.8-0.8 being located at the far-right-end of the tail (see the vertical dash-dotted line in Fig. 11b). Same conclusion can be drawn for the along-wind sections as well (not shown). Therefore, the exponent 0.8-0.8 is statistically significant with respect to a pure white noise. This is likely since our analyses focused on periods with high temperature variance and hence on periods with high signal-to-noise ratio.
Figure 12: The ensemble-averaged DTS temperature spectra are shown for (a) temporal and (b) spatial domains, respectively. The error bars indicate one standard deviation with respect to the ensemble mean. Only along-wind DTS cable sections (sections 1-5) are used for illustration purposes. The red and black dash dotted lines in (a) indicate the 1-1 and 5/3-5/3 spectral slopes, respectively. The magenta dash-dotted line in (b) indicates a specific slope of $\approx-1.13$ (see Eq. 19).
The second type of noise is related to spectral features of DTS data (cheng2017failure). For the along-wind DTS cable sections, in Fig. 12, we show the spectra of temperature fluctuations computed over the temporal (Fig. 12a) and spatial domains (Fig. 12b), respectively. The error bars indicate one standard deviation with respect to the ensemble mean and the varying shades of gray represent different sections. The spectral amplitudes ($S_{TT}$) are normalized by the temperature variances ($\sigma^{2}_{T}$) estimated from the EC measurements and the frequencies are scaled with the standard surface layer scaling ($2\pi fz/U_{C}$) of kaimal1972spectral. From Fig. 12a, one could see that the DTS temperature spectra in the temporal domain follow a 5/3-5/3 slope at smaller scales while displaying a 1-1 slope at larger scales. Upon comparison of DTS temporal spectra with the along-wind EC temperature spectra, a satisfactory agreement is found between the two (see Fig. S5 in the Supplementary Material). However, since the temporal resolution of DTS data ($2.3\ \rm s$) is lower than the EC one ($0.1\ \rm s$), it does not fully capture the location of the spectral peak. On the other hand, for the spatial domain, the DTS temperature spectra show a rather narrow region of 5/3-5/3 slope (Fig. 12b). This is followed by a region with a spectral slope larger than 5/3-5/3. Precisely, this enhancement is equal to $5/3+(5/3-1)\times 0.8\approx-1.13$ (see Eq. 19).
By following cheng2017failure, we decided to truncate this enhanced region from the spatial DTS spectrum by applying a smoothing filter and reevaluating the phase-dependent convection speeds. This decision was undertaken because upon a closer look at Fig. 5b, it appears that the presence of a 1/3-1/3 slope could be there up to the wavenumber range $\kappa z<3$. We wanted to confirm whether the scaling of scale-dependent convection speeds would revert to ${(\kappa z)}^{-1/3}$ by removing this enhanced region from the spatial DTS spectrum. From Fig. 12b, one could see that the enhancement in spectral powers occurs at scaled wavenumbers of $\kappa z\geq 3$. The limit $\kappa z=3$ is nearly equivalent to a wavelength of 6 m, which corresponds to a spatial separation of 3 m. Therefore, to truncate the spectra beyond $\kappa z\geq 3$, one could average 10 consecutive spatial points, which artificially would reduce the spatial resolution from 30 cm to 3 m. For this artificially reduced data, we computed the spatial DTS spectrum and as expected, the spectrum of this smoothed dataset only revealed the 5/3-5/3 spectral slope (see Fig. S6 in the Supplementary Material). Next, we calculated the scale-dependent convection speeds based on this filtered data to see if indeed those would revert to the 1/3-1/3 scaling of cheng2017failure. This does not appear to be the case, as even for this filtered data, the visual inspection of Fig. S7 in the Supplementary Material suggests that the convection speeds have a preference towards a 0.8-0.8 scaling for all the five cable sections. To confirm this visual assessment, corresponding to the wavenumber range $\kappa z\geq 1$, the exponent was computed by doing a linear regression between $\log[U_{C}(\kappa z)/U_{C}]$ and $\log[\kappa z]$. The answer turned out to be nearly 0.8-0.8 accompanied by an $R^{2}$ value larger than 0.9.
The third type of noise could arise due to a linear-interpolation operation to remove the temperature values at specific locations of the wooden structures. Though the number of points being removed were only 6% of the total cable length, we artificially increased the missing data to be as large as 20% of the cable length. Thereafter, we replaced the missing values with linear interpolation and computed the $U_{C}(\kappa z)$ curves. As shown in Fig. S8 of the Supplementary Material, the results appeared to be insensitive to the linear-interpolation operation. Specifically, with respect to the original signal, one could not detect any appreciable change in the estimations of $U_{C}(\kappa z)$ and therefore, the 0.8-0.8 scaling survived.